OpenStudy (anonymous):

Find the critical number of the function. g(t) = | 3t - 5 | will fan and medal!!!

OpenStudy (aum):

\[ 3t - 5 = 0, \text{when }~t = \frac 53\\ \text{For }~~t \gt \frac 53, ~~g(t) = 3t - 5;~~~~~g'(t) = 3 \\ \text{For }~~t \lt \frac 53, ~~g(t) = -3t + 5;~~g'(t) = -3 \\ \text{At }~~t = \frac 53, ~~g'\left (\frac 53^+\right ) = 3; ~~ g'\left (\frac 53^-\right ) = -3 \\ \text{Therefore, the derivative does not exist at }t = \frac 53. \\ t = \frac 53\text{ is a critical point.} \]